# Deep Improvement Supervision

### v2

@@ Abstract
[…]
Machine Learning, ICML
@@ 1 Introduction
The goal of reasoning models is to start from a set of specific examples or observations and infer a general rule or pattern. Such models systematically manipulate and connect pieces of information to infer new conclusions and solve problems not explicitly stated in the training data. Building such models is a key challenge for current machine learning approaches. The use of iterative refinement loops on a model’s outputs has become a major driver of progress in deep learning reasoning, a trend evident in both small and large-scale architectures.
At a high level, the reasoning process in large language models (LLMs) can be viewed as a product of this same principle: a recursive, loop-like refinement of the model’s own outputs (hurst2024gpt; jaech2024openai). Their primary reasoning mechanism, Chain-of-Thought (CoT) prompting (wei2022chain), relies on externalizing reasoning into a sequence of generated text steps. However, this process in LLMs demands extensive computational infrastructure, making it highly expensive.
Figure 1: Blueprint of the discrete diffusion process on the ARC. Starting from the input x, following timestep $t$, we generate diffusion steps to the target y (chollet2019measure).
For smaller models, this idea is also proved itself, looped transformers (giannou2023looped; yang2023looped) repeatedly apply the same transformer block with input injection at each step and achieve better performance than a standard one forward pass transformer on reasoning and meta learning tasks, why using 10x less number of parameters (yang2023looped).
Recently, models like Hierarchical Reasoning Models (HRM) (wang2025hierarchical) and Tiny Recursive Models (TRM) (jolicoeur2025less), was built on the similar idea of reusing model output repeatedly in both the input and latent spaces. These models have demonstrated very impressive performance and even outperformed multi-billion parameter LLM models on complicated ARC-AGI (chollet2019measure) tasks. However, a path of building reasoning models based on looped inference introduces a profound question. When a reasoning model executes a refinement step, what guaranties that it is on a path to a better answer? How can we guide the model to ensure that each step is verifiably closer to the right conclusion?
In this paper, we propose a learning algorithm that makes the reasoning process more task-oriented. Our main finding is that guiding the model to predict intermediate steps via discrete diffusion (see Fig. 1) during iterative reasoning significantly reduces training complexity and improves generalization. We demonstrate that, with a target for each supervision step, the no-grad cycles proposed for TRM and HRM can be avoided. Additionally, by performing a fixed number of refinement steps, we simplify training because the need to halt is eliminated.
To formalize our approach and hypotheses regarding the efficacy of TRMs, we first model the latent reasoning process through distinct theoretical lenses, establishing a tractable and well-defined formulation. Specifically, we show that a TRM can be interpreted as a form of classifier-free diffusion guidance (Sec. 3.1). Furthermore, we demonstrate how asymmetric latent reasoning facilitates policy improvement that exceeds the policy inherent in the training data (frans2025diffusion) (Sec. 3.2). Finally, based on this formulation, we theoretically justify the benefits of step-wise guidance.
@@ 2.1 Hierarchical Reasoning Models
A looped (giannou2023looped) and universal (dehghani2018universal) transformers repeatedly applies the same transformer block, with input injection each time, and is trained to make its intermediate loop output correct. This model was applied for the various tasks, showing the improvement over the single step models (yang2023looped). Based on this idea, Hierarchical Reasoning Models (HRMs) (wang2025hierarchical) are supervised sequence-to-sequence models that perform recursive refinement of a prediction by interleaving two small recurrent networks that operate at different update frequencies. Let $\tilde{\mathbf{x}}\in\mathcal{V}^{L}$ denote an input sequence of length $L$ on a vocabulary $\mathcal{V}$, and let $\mathbf{y}\in\mathcal{V}^{L}$ be the desired output. HRM uses an input embedding $f_{I}$, two recurrent reasoning modules a low-level module $f_{L}$ and a high-level module $f_{H}$ and an output head $f_{O}$. After embedding $\mathbf{x}=f_{I}(\tilde{\mathbf{x}})\in\mathbb{R}^{L\times D}$, HRM carries two latent states $\mathbf{z}_{L},\mathbf{z}_{H}\in\mathbb{R}^{L\times D}$ through the supervision steps. Within a forward pass, it performs $n$ updates of $f^{\phi}_{L}$ for every update of $f^{\psi}_{H}$ and repeats this $T$ times before decoding with $f_{O}$. A typical schedule used in previous work is $n=2$ and $T=2$. The final prediction is $\hat{\mathbf{y}}=\operatorname*{arg\,max}f_{O}(\mathbf{z}_{H})$. During a forward pass, HRM evaluates the following updates:
@@
Adaptive Computational Time. Training-time efficiency is improved by a learned halting mechanism (ACT). A small head predicts whether to stop iterating on the current example or continue; the published implementation trains this with a halting loss and an additional continue loss that requires an extra forward pass, effectively doubling forward compute per optimization step (graves2016act). The test-time evaluation runs a fixed maximum number of supervision steps to maximize accuracy.
@@ 2.2 Tiny Recursive Models
Tiny Recursive Models (jolicoeur2025less) retain the core idea of iterative refinement but collapse HRM’s complexity into a single tiny network and a simpler recursion scheme. In the TRM setup, $\mathbf{z}_{H}$ is the state that the model reads out to produce the answer (the output head is applied to $\mathbf{z}_{H}$: $\hat{\mathbf{y}}=\arg\max f_{O}(\mathbf{z}_{H})$). $\mathbf{z}_{L}$ is a working memory state that’s updated using the input $\mathbf{x}$ and the current answer, and is then used to update $\mathbf{z}_{H}$. Because the loss is applied on the prediction from $\mathbf{z}_{H}$, optimization pressure makes $\mathbf{z}_{H}$ look like (encode) the current solution. On the other hand, $\mathbf{z}_{L}$ is only supervised indirectly through its effect on $\mathbf{z}_{H}$, so it is free to be an internal reasoning representation rather than a decodable solution. Within same $f^{\phi}$ network for $L$ and $H$ module, TRM repeats each recursion equal to HRM Eq. 1.
@@
The prediction is made from $\mathbf{z}_{H}$ through the output head and trained with cross-entropy. This asymmetry (only $\mathbf{z}_{H}$ sees final $\mathbf{z}^{n+1}_{L}$; and only $\mathbf{z}_{H}$ is decoded and penalized) naturally pushes $\mathbf{z}_{H}$ towards the space of valid answers, while $\mathbf{z}_{L}$ becomes the latent reasoning scratchpad that helps improve the next $\mathbf{z}_{H}$. The TRM paper explicitly reframes this: $\mathbf{z}_{H}$ is simply the current (embedded) solution… $\mathbf{z}_{L}$ is a latent feature that does not directly correspond to a solution but can be transformed into one by $f_{H}$. It was show on the Sudoku example that when you reverse-embed + argmax, the tokenized $\mathbf{z}_{H}$ looks like the solved grid, while tokenized $\mathbf{z}_{L}$ is not realted to the solution.
@@
From the notation perspective, we want to note that TRM renames $\mathbf{z}_{H}$ to $\mathbf{y}$ (the running answer) and $\mathbf{z}_{L}$ to $\mathbf{z}$ (latent reasoning). The loop becomes: update $\mathbf{z}$ using $(\mathbf{x},\mathbf{y},\mathbf{z})$; then update $\mathbf{y}$ using $(\mathbf{y},\mathbf{z})$. Carrying both across deep-supervision steps lets the model iterate: $\mathbf{z}$ remembers how it got to the current guess (like a chain-of-thought), and $\mathbf{y}$ stores the current guess itself. TRM trains a single halting probability via binary cross-entropy against correctness. Let $\mathcal{L}_{\text{task}}=\mathrm{CE}(f_{O}(\mathbf{y}),\mathbf{y}_{\text{true}})$ be the prediction cross-entropy loss and $\mathcal{L}_{\text{halt}}=\mathrm{BCE}(q(\mathbf{y}),\hat{\mathbf{y}}=\mathbf{y}_{\text{true}})$ the halting loss with $q(\cdot)$ a scalar head. An optimization step iterates up to $N_{\text{sup}}$ supervision steps, performing T1T-1 no-grad recursion cycles, then one with gradients, detaching $(\mathbf{y},\mathbf{z})$ between supervision steps. Early stopping within a minibatch is permitted by using the halting signal.
2.3 Diffusion Guidance as Policy Improvement
We use the policy improvement point of view to describe the latent reasoning process of TRM. For clarity, we provide a short background section below.
@@ 2.3 Diffusion Guidance as Policy Improvement
Recent work formalizes a tight connection between classifier-free guidance (CFG) in diffusion/flow models and policy improvement in reinforcement learning (RL) (frans2025diffusion). This work provides a framework for the analysis of diffusion models as RL methods. For a state $\mathbf{s}$ and an action/output $\mathbf{y}$, the core construction, termed classifier-free guidance RL (CFGRL), parameterizes a target policy as:
$\pi(y\mid\mathbf{s})\;\propto\;\hat{\pi}(y\mid\mathbf{s})\;f\!\big(A_{\hat{\pi}}(\mathbf{s},\mathbf{y})\big),$ (2)
where $f:\mathbb{R}\!\to\!\mathbb{R}_{\geq 0}$ is a nonnegative monotonically increasing function of the advantage $A^{\pi}(\mathbf{s},\mathbf{y})=Q^{\pi}(\mathbf{s},\mathbf{y})-V^{\pi}(\mathbf{s})$ (sutton1998reinforcement), and $\hat{\pi}$ is a reference policy. According to the formal proof of the authors in the appendix (Theorem 1) (frans2025diffusion), a new policy generated from a carefully weighted combination of previous policies is guaranteed to achieve a higher expected return than the original reference policy. This result generalizes earlier findings in bandits and provides a pathway for improving policies.
We can say that the reference policy is learned from given data, or we can say that this is any non-optimal policy that we want to improve. Following (2) $\pi$ provably improves over $\hat{\pi}$; moreover, attenuating the optimality factor with an exponent w{w} yields a family $\pi_{w}\propto\hat{\pi}f(A)^{w}$ whose expected return increases with w{w}. This recovers KL-regularized policy improvement as a special case, where the optimal solution has the same product form with $f(A)=\exp(A/\beta)$, tying w{w} to the trust-region strength $\beta$ without retraining.
Crucially, CFGRL shows that diffusion guidance composes these factors at the score level. Let $\mathbf{o}\in\{\emptyset,1\}$ indicate optimality, defined by $f$ as $p(\mathbf{o}|\mathbf{s},\mathbf{y})=f(A(s,\mathbf{y}))$, then the product policy of Eq. (2) can now be equivalently defined as: $\pi(y\mid\mathbf{s})\;\propto\;\hat{\pi}(y\mid\mathbf{s})p(\mathbf{o}|\mathbf{s},\mathbf{y})$.
As such, the score of the product policy above can be represented as the sum of two factors $\nabla_{a}\log\pi(y|\mathbf{s})=\nabla_{a}\log\hat{\pi}(y|\mathbf{s})+\nabla_{a}\log p(\mathbf{o}|\mathbf{s},y)$. Then, using Bayes rule and w{w}, CFG produces an score:
$\nabla_{y}\log\pi_{w}(y\mid\mathbf{s})\;=\;\nabla_{y}\log\hat{\pi}(y\mid\mathbf{s})+\;$
${w}\Big(\nabla_{y}\log\hat{\pi}(y\mid\mathbf{s},\mathbf{o}{=}1)-\nabla_{y}\log\hat{\pi}(y\mid\mathbf{s})\Big),$ (3)
which corresponds to sampling from $\pi_{w}\!\propto\!\hat{\pi}(y\mid\mathbf{s})\,p(\mathbf{o}{=}1\mid s,\mathbf{y})^{w}$. Please see (frans2025diffusion) for other derivation details. This realizes controllable policy improvement at test time by tuning w{w}, while training remains a simple supervised diffusion/flow objective with classifier-free dropout of $\mathbf{o}$. Overall, CFGRL preserves the stability and simplicity of generative supervised training while providing a principled, test-time knob for step-size in policy improvement. Empirically, CFGRL shows policy improvement and yields consistently better returns than the reference policy $\hat{\pi}$
3 Different Faces of Latent Reasoning
In this section, we decompose the architectural elements of the TRM to demonstrate its relationship with well‑established frameworks in generative modeling and reinforcement learning. The goal is to provide a formal hypothesis for why TRM works by connecting the model’s asymmetric latent updates to Classifier‑Free Guidance (CFG) and to Implicit Policy Improvement.

### v3

# Your Latent Reasoning is Secretly Policy Improvement Operator

@@ Abstract
Recently, small models with latent recursion have obtained promising results on complex reasoning tasks. These results are typically explained by the theory that such recursion increases a network’s depth, allowing it to compactly emulate the capacity of larger models. However, the performance of recursively added layers remains behind the capabilities of one‑pass models with the same feed-forward depth. This means that in the looped version, not every recursive step effectively contributes to depth. This raises the question: when and why does latent reasoning improve performance, and when does it result in dead compute? In our work, we analyze the algorithms that latent reasoning provides answer to this question. We show that latent reasoning can be formalized as a classifier‑free guidance and policy improvement algorithm. Building on these insights, we propose to use a training schemes from reinforcement learning and diffusion methods for latent reasoning models. Using the Tiny Recursive Model as our testbed, we show that with our modifications we can avoid dead compute steps and reduce the total number of forward passes by 18× while maintaining performance. Broadly speaking, we show how a policy improvement perspective on recursive steps can explain model behavior and provide insights for further improvements.
Reasoning, Latent, TRM, RL, ARC-AGI, Diffusion
@@ 1 Introduction
The goal of reasoning models is to start from a set of specific examples or observations and infer a general rule or pattern. Such models systematically manipulate and connect pieces of information to infer new conclusions and solve problems not explicitly stated in the training data.
Recently, the use of latent reasoning and iterative refinement loops has become a major driver of progress in deep learning reasoning. Looped transformers (Giannou et al., 2023; Yang et al., 2023) repeatedly apply the same transformer block with input injection at each step and achieve better performance than a standard one forward pass transformer on reasoning and meta learning tasks, why using 10x less number of parameters (Yang et al., 2023).
Recently, models like Hierarchical Reasoning Models (HRM) (Wang et al., 2025) and Tiny Recursive Models (Jolicoeur-Martineau, 2025), were built on the similar idea of reusing model output repeatedly in both the input and latent spaces. These models have demonstrated impressive performance and even outperformed multi-billion parameter LLM models on complicated ARC-AGI (Chollet, 2019).
In our paper we ask a question when an why is a simulated effective layer really effective? And after answering this, we want to know a better path on how such a model can be improved. This latent reasoning does some algorithm using the steps in latent space, but what kind of algorithm is actually running in this latent space. In our paper, we are not seeking for some interoperability of the latent representation, our goal is to formally bridge the algorithm of latent space reasoning to the well-known frameworks.
For this, we demonstrate how latent reasoning facilitates the policy improvement algorithm that aims to improve an advantage margin between the one-step output and the recurrence one (Frans et al., 2025). based on our findings, we propose an improved algorithm that makes the reasoning process more task-oriented.
TRMs are difficult to train efficiently: supervision is primarily applied to the final output, leaving early refinement steps with weak credit assignment and making optimization brittle in the small-data regime. We propose Deep Improvement Supervision (DIS), a training method that supplies step-wise intermediate targets during recursion. In our main implementation, these targets are generated by a monotonic discrete corruption schedule of the ground-truth output, turning each recursion step into a supervised subgoal.
@@ 2.1 Hierarchical Reasoning Models
A looped (Giannou et al., 2023) and universal (Dehghani et al., 2018) transformer repeatedly applies the same transformer block, with input injection each time, and is trained to make its intermediate loop output correct. This model was applied for various tasks, showing an improvement over the single step models (Yang et al., 2023). Based on this idea, Hierarchical Reasoning Models (HRMs) (Wang et al., 2025) are supervised sequence-to-sequence models that perform recursive refinement of a prediction by interleaving two small recurrent networks that operate at different update frequencies. Let $\tilde{\mathbf{x}}\in\mathcal{V}^{L}$ denote an input sequence of length $L$ in a vocabulary $\mathcal{V}$, and let $\mathbf{y}\in\mathcal{V}^{L}$ be the desired output. HRM uses an input embedding $f_{I}$, two recurrent reasoning modules, a low-level module $f_{L}$ and a high-level module $f_{H}$ and an output head $f_{O}$.
After embedding $\mathbf{x}=f_{I}(\tilde{\mathbf{x}})\in\mathbb{R}^{L\times D}$, HRM carries two latent states $\mathbf{z}_{L},\mathbf{z}_{H}\in\mathbb{R}^{L\times D}$ through the supervision steps. Within a forward pass, it performs $n$ updates of $f^{\phi}_{L}$ for every update of $f^{\psi}_{H}$ and repeats this $T$ time before decoding with $f_{O}$. A typical schedule used in previous work is $n=2$ and $T=2$. The final prediction is $\hat{\mathbf{y}}=\operatorname*{arg\,max}f_{O}(\mathbf{z}_{H})$. During a forward pass, HRM evaluates the following updates:
@@
Adaptive Computational Time. Training-time efficiency is improved by a learned halting mechanism (ACT). A small head predicts whether to stop iterating on the current example or continue; the published implementation trains this with a halting loss and an additional continue loss that requires an additional forward pass, effectively doubling forward compute per optimization step (Graves, 2016). The test-time evaluation runs a fixed maximum number of supervision steps to maximize accuracy.
@@ 2.2 Tiny Recursive Models
Tiny Recursive Models (Jolicoeur-Martineau, 2025) retain the core idea of iterative refinement but collapse HRM’s complexity into a single tiny network and a simpler recursion scheme. In the TRM setup, $\mathbf{z}_{H}$ is the state that the model reads out to produce the answer (the output head is applied to $\mathbf{z}_{H}$: $\hat{\mathbf{y}}=\arg\max f_{O}(\mathbf{z}_{H})$). $\mathbf{z}_{L}$ is a working memory state that is updated using the input $\mathbf{x}$ and the current answer, and is then used to update $\mathbf{z}_{H}$. Because the loss is applied on the prediction from $\mathbf{z}_{H}$, optimization pressure makes $\mathbf{z}_{H}$ look like (encode) the current solution. On the other hand, $\mathbf{z}_{L}$ is only indirectly supervised through its effect on $\mathbf{z}_{H}$, so it is free to be an internal reasoning representation rather than a decodable solution. Within same $f^{\phi}$ network for $L$ and $H$ module, TRM repeats each recursion equal to HRM Eq. 1.
@@
The prediction is made from $\mathbf{z}_{H}$ through the output head and trained with cross-entropy. This asymmetry (only $\mathbf{z}_{H}$ sees final $\mathbf{z}^{n+1}_{L}$; and only $\mathbf{z}_{H}$ is decoded and penalized) naturally pushes $\mathbf{z}_{H}$ towards the space of valid answers, while $\mathbf{z}_{L}$ becomes the latent reasoning scratchpad that helps improve the next $\mathbf{z}_{H}$.
The TRM paper explicitly reframes this: $\mathbf{z}_{H}$ is simply the current (embedded) solution… $\mathbf{z}_{L}$ is a latent feature that does not directly correspond to a solution but can be transformed into one by $f_{H}$”. It was shown on the Sudoku example that when you reverse-embed + argmax, the tokenized $\mathbf{z}_{H}$ looks like the solved grid, while tokenized $\mathbf{z}_{L}$ is not realted to the solution.
@@
From the notation perspective, we want to note that TRM renames $\mathbf{z}_{H}$ to $\mathbf{y}$ (the running answer) and $\mathbf{z}_{L}$ to $\mathbf{z}$ (latent reasoning). The loop becomes: update $\mathbf{z}$ using $(\mathbf{x},\mathbf{y},\mathbf{z})$; then update $\mathbf{y}$ using $(\mathbf{y},\mathbf{z})$. Carrying both across deep-supervision steps lets the model iterate: $\mathbf{z}$ remembers how it got to the current guess (like a chain-of-thought), and $\mathbf{y}$ stores the current guess itself. TRM trains a single halting probability via binary cross-entropy against correctness.
Let $\mathcal{L}_{\text{task}}=\mathrm{CE}(f_{O}(\mathbf{y}),\mathbf{y}_{\text{true}})$ be the prediction cross-entropy loss and $\mathcal{L}_{\text{halt}}=\mathrm{BCE}(q(\mathbf{y}),\hat{\mathbf{y}}=\mathbf{y}_{\text{true}})$ the halting loss with $q(\cdot)$ a scalar head. An optimization step iterates up to $N_{\text{sup}}$ supervision steps, performing T1T-1 no-grad recursion cycles, then one with gradients, detaching $(\mathbf{y},\mathbf{z})$ between supervision steps. Early stopping within a minibatch is permitted by using the halting signal.
2.3 Policy Improvement Algorithm
We consider a Markov decision process with state space $\mathcal{S}$, action space $\mathcal{A}$, and a stochastic policy $\pi(a\mid s)$. Given a reference policy $\hat{\pi}$, we define the discounted state-value and action-value functions as $V^{\hat{\pi}}(s)$ and $Q^{\hat{\pi}}(s,a)$, respectively. The advantage function is defined as
$A^{\hat{\pi}}(s,a)=Q^{\hat{\pi}}(s,a)-V^{\hat{\pi}}(s).$
[…]
$\tilde{J}(\pi)=\mathbb{E}_{s\sim p_{\hat{\pi}}(s)}\left[\mathbb{E}_{a\sim\pi(a\mid s)}\left[A^{\hat{\pi}}(s,a)\right]\right],$ (2)
where $p_{\pi}(s)$ denotes the discounted visitation distribution under $\pi$ (Sutton et al., 1998). To control deviations from the policy, it is common to introduce a KL regularization:
$J(\pi)$ $=\mathbb{E}_{\tau\sim p(\tau\mid\pi)}\left[\sum_{t}\gamma^{t}r(s_{t},a_{t})\right]$
$-\beta\,\mathbb{E}_{s\sim p_{\pi}(s)}\left[D_{\mathrm{KL}}\!\left(\pi(\cdot\mid s)\,\|\,\hat{\pi}(\cdot\mid s)\right)\right],$ (3)
where $\beta>0$ controls the tradeoff between reward maximization and adherence to $\hat{\pi}$. The solution to the KL-regularized policy improvement problem in Equation (3) takes the form of a product policy:
$\pi(a\mid s)\propto\hat{\pi}(a\mid s)\exp\!\left(\tfrac{1}{\beta}A^{\hat{\pi}}(s,a)\right),$ (4)
which corresponds to exponentiated advantage reweighting. This shows that policy improvement can be achieved by multiplicatively reweighting a reference policy using any monotonic function of advantage, providing a unifying view of trust-region methods, advantage-weighted regression, and KL-regularized reinforcement learning objectives.
@@ 2.3 Policy Improvement Algorithm
Recently was shown that policy improvement framework does not necessarily rely on explicitly learning a value function. Recent work formalizes a tight connection between classifier-free guidance (CFG) in diffusion/flow models and policy improvement in reinforcement learning (Frans et al., 2025). This work provides a framework for the analysis of diffusion models as RL methods. For a state $\mathbf{s}$ and an action/output $\mathbf{a}$, the classifier-free guidance RL (CFGRL) method parameterizes a target policy as:
$\pi(a\mid\mathbf{s})\;\propto\;\hat{\pi}(a\mid\mathbf{s})\;f\!\big(A_{\hat{\pi}}(\mathbf{s},\mathbf{a})\big),$ (5)
where $f:\mathbb{R}\!\to\!\mathbb{R}_{\geq 0}$ is a nonnegative monotonically increasing function of the advantage $A^{\pi}(\mathbf{s},\mathbf{a})=Q^{\pi}(\mathbf{s},\mathbf{a})-V^{\pi}(\mathbf{s})$ (Sutton et al., 1998), and $\hat{\pi}$ is a reference policy.
3 Formalization of Latent Reasoning as Policy Improvement
Both (TRMs) and (HRMs) are often motivated by effective depth: reusing a small parameter-shared network is said to emulate very deep computation with few parameters. In HRMs this story is tied to Deep Equilibrium (DEQ) models, where one assumes convergence to a fixed point and invokes the Implicit Function Theorem to justify differentiating through only the last iterate (Bai et al., 2019). However, follow-up analyses indicate that the fixed-point assumption is typically violated in practice: latent residuals remain non-negligible under truncation, and a one-step DEQ approximation does not fully explain the empirical gains (Jolicoeur-Martineau, 2025).
This paper adopts a different organizing principle. Rather than asking whether recursion converges, we ask: what does one latent-reasoning step do to the model’s output distribution? Our central claim is that a single TRM step implements a policy improvement update: it takes a reference policy (the model’s current guess) and transforms it into an improved policy using an internally computed signal.
Notation. At recursion step $t$ we write the TRM state as
$s_{t}\;=\;(x,z_{L}^{t},z_{H}^{t}),$